# Bias in Large Language Models: Origin, Evaluation, and Mitigation

### v1

@@ A.2.4 Recommender Systems
Recommender Systems in the context of LLMs are designed to predict and suggest personalized content or items of interest to users based on their preferences, behavior, or contextual information. These systems analyze patterns in user data, such as past interactions, preferences, or demographic information, to provide personalized recommendations, such as suggesting movies, products, books, or articles that the user might enjoy or find useful (Zhang et al., 2019a; Hua et al., 2023).
@@
    • Product Recommendations Based on Gender Stereotypes: A recommender system might suggest beauty products, fashion items, or household goods primarily to women while recommending electronics, tools, or sports equipment predominantly to men. This bias reinforces traditional gender roles and can lead to irrelevant or unappealing recommendations for users whose interests do not align with these stereotypes (Ekstrand et al., 2018).
@@ A.2.5 Machine Translation
    • Gendered Language Mismatch: When translating from a gender-neutral language like Turkish or Finnish into a gendered language like English or Spanish, the model might introduce gender bias by assigning gendered pronouns or roles based on stereotypes. For example, the Turkish sentence “O bir doktor” (which means “He/She is a doctor”) might be translated into English as “He is a doctor,” reflecting the stereotype that doctors are male (Prates et al., 2020).
@@
    • Translation of Age-Related Idioms: When translating age-related idioms or expressions, a biased translation system might reinforce negative stereotypes. For instance, translating a phrase like “old people are slow” into another language might retain the negative connotation or even intensify it if the target language has a stronger cultural bias against the elderly. (Hovy and Søgaard, 2015).
@@
    • Bias Toward Dominant Cultures: A machine translation system might favor translations that align with Western cultural norms over those of less dominant cultures. For instance, translating phrases related to food, clothing, or customs might reflect Western standards, even when the source text belongs to a non-Western culture. An example could be translating a traditional Chinese clothing item, “Qipao,” simply as “dress,” which dilutes the cultural significance (Sennrich, 2016).

### v2

@@ A.2.4 Recommender Systems
Recommender systems in the context of LLMs are designed to predict and suggest personalized content or items of interest to users based on their preferences, behavior, or contextual information. These systems analyze patterns in user data, such as past interactions, preferences, or demographic information, to provide personalized recommendations, such as suggesting movies, products, books, or articles that the user might enjoy or find useful (Zhang et al., 2019; Hua et al., 2023).
@@
    • Product recommendations based on gender stereotypes: A recommender system might suggest beauty products, fashion items, or household goods primarily to women, while recommending electronics, tools, or sports equipment predominantly to men. This bias reinforces traditional gender roles and can lead to irrelevant or unappealing recommendations for users whose interests do not align with these stereotypes (Ekstrand et al., 2018).
@@ A.2.5 Machine Translation
    • Gendered language mismatch: When translating from a gender-neutral language like Turkish or Finnish into a gendered language like English or Spanish, the model might introduce gender bias by assigning gendered pronouns or roles based on stereotypes. For example, the Turkish sentence “O bir doktor” (which means “He/She is a doctor”) might be translated into English as “He is a doctor”, reflecting the stereotype that doctors are male (Prates et al., 2020).
@@
    • Translation of age-related idioms: When translating age-related idioms or expressions, a biased translation system might reinforce negative stereotypes. For instance, translating a phrase like “old people are slow” into another language might retain the negative connotation or even intensify it if the target language has a stronger cultural bias against the elderly (Hovy and Søgaard, 2015).
@@
    • Bias toward dominant cultures: A machine translation system might favor translations that align with Western cultural norms over those of less dominant cultures. For instance, translating phrases related to food, clothing, or customs might reflect Western standards, even when the source text belongs to a non-Western culture. An example could be translating a traditional Chinese clothing item, “Qipao”, simply as “dress”, which dilutes the cultural significance (Sennrich, 2016).